c. Now assume a transmit power of 250 W at the earth station. What is the power received at the satellite antenna in dBW and in W?
Answer #1

Note that

Pr =Pt/(4 π d2)

where Pr = received power, Pt = transmitted power and d is distance from transmitter in meters.

Pr = 250 / (4 x 3.14 x d2) W

Note that you have not provided the distance between the earth and the satellite

Pr = 19.9 / d2 W

Power in dbW is given by -

The power in decibel-watts (P(dBW) is equal to 10 times base 10 logarithm of the power in PdBW) = 10 . log10(PW) / 1W)

Put the value of Pr in place of P(W) to obtain P(dbW)

I have tried to explain it in very simple language and I hope that i have answered your question satisfactorily.Leave doubts in comment section if any.